Tom throws a ball into the air. The ball travels on a parabolic path represented by the equation h = -8t^2 + 40t, where h represents the height of the ball above the ground and t represents the time in seconds.
How many seconds does it take the ball to reach its highest point?
What ordered pair represents the highest point that the ball reaches as it travels through the air?

Answer by jorel1380(3719)

The highest the ball will reach, given the quadratic equation ax˛+bx+c to describe its' motion, will be represented by -b/2a. This is the vertex of the parabola described in the equation. So:
-(40/2(-8)=40/16=2.5 seconds before the ball reaches its' highest point.
The ordered pair that represents this is:
(2.5, (-8(2.5˛))+40(2.5), or
(2.5,50) ☺☺☺☺